OpenStudy (anonymous):

which point is NOT a solution of x+2y≤ 8? a. (0,4) b.(-2,3) c.(4,4) d. (6,0)

OpenStudy (anonymous):

please help

OpenStudy (ipwnbunnies):

The caveman way of doing this is to plug each choice in. Does it satisfy the equation?

OpenStudy (ipwnbunnies):

For A, plug in 0 for x and 4 for y. Is it less than or equal to 8? 0 + 2(4) =< 8? 8 = 8, so the point (0,4) is a solution.

OpenStudy (anonymous):

Thanks and so that all you bascially dojust plug in the points to see if they are correct or not ..

OpenStudy (ipwnbunnies):

Right.

OpenStudy (anonymous):

do you mind helping me with another one?

OpenStudy (ipwnbunnies):

Is it the same concept?

OpenStudy (anonymous):

yes but it isn't in y=mx+b form

OpenStudy (ipwnbunnies):

Then put it in mx + b form?

OpenStudy (the_fizicx99):

Simply plug in those coordinates and see which is not true, using A, \(\ \sf \Large (\color{orange}{x},\color{blue}{y}) \) So, \(\ \sf \Large (\color{orange}{0},\color{blue}{4}) \) \(\ \sf \Large \color{orange}{x} + 2\color{blue}{y} \le 8 \longrightarrow \color{orange}{0} + 2(\color{blue}{4}) \le 8 \longrightarrow 0 + 8 \le 8\) \(\ \sf \Large Is ~~ 8 \le 8 ? ~~ Yes, 8 ~~ is ~~ equal ~~ to ~~ 8! \) So A doesn't work.

OpenStudy (anonymous):

y<-3 A.(1,-4) B.(3,0) C.(-2,-5) D.(8,-7) you see it is the exact same concept, but i dont understand how you would plug in the numbers to this one..

OpenStudy (ipwnbunnies):

Well, this one is easier. We want the point that lies underneath the horizontal line, y = 3.

OpenStudy (anonymous):

@tHe_FiZiCx99 then how do i find the right way, can you help me also..?

OpenStudy (anonymous):

i see @iPwnBunnies so how would this help me find the one that doesn't work instead of the one that does work?

OpenStudy (ipwnbunnies):

y = -3**

OpenStudy (the_fizicx99):

Sure, sec

OpenStudy (anonymous):

@iPwnBunnies so this would indicate that the first one does work, but what about the ones that don't, because its asking me how to find the one that doesn't work

OpenStudy (ipwnbunnies):

No, the first one works. The point is (1, -4). The y-value is -4. Isn't that less than -3?

OpenStudy (anonymous):

@iPwnBunnies yes it does work, but i need to find the one that doesn't work you see that the problem, i need to find the one that DOESN'T work.

OpenStudy (ipwnbunnies):

Well, in fact, you didn't write instructions for the second problem lol.

OpenStudy (anonymous):

ohh ahah im sorry i thought we were talking about the first one ah my bad Cx

OpenStudy (ipwnbunnies):

Since y < -3 has to be satisfied, a point with a Y-VALUE OVER -3 won't be a solution to the inequality.

OpenStudy (the_fizicx99):

So as Pwn said, this is a bit easier! Simply plot a horizontal line at (0,-3) and it has to be dashed. Also, it's \(\ \sf \Large Less ~ than \) so you'll shade below the line. Now look at option A, is (1, -4) -4 is indeed less than -3

OpenStudy (anonymous):

@iPwnBunnies what do you mean? and the first one is the same thing i need to find the one that doesn't work also

OpenStudy (anonymous):

@tHe_FiZiCx99 well can you help me on number one

OpenStudy (the_fizicx99):

Lopez, all 3 options work actually. Is it asking for a coordinate pair that Doesn't fit y < -3 ?

OpenStudy (anonymous):

yes its asking for the ordered pair that isn't a solution of y<-3..

OpenStudy (the_fizicx99):

I just read your replies and it does seem you're looking for a par that doesn't satisfy y <-3 Well look at it this way! Is -4 less than -3? Is 0 less than -3? Is -5 less than -3? Is -7 less than -3? One of them isn't less than! Once you answer them, you'll see which option it is :>

OpenStudy (the_fizicx99):

pair*

OpenStudy (anonymous):

so the option so it would be D.?

OpenStudy (the_fizicx99):

Hm, is -7 LESS THAN -3 ?

OpenStudy (anonymous):

yes it is..

OpenStudy (anonymous):

isn't it?

OpenStudy (the_fizicx99):

So D doesn't work! You're looking for one that s greater than -3, one that doesn't go into the shaded area.

OpenStudy (anonymous):

B. would be thr right answer than i thought we werte looking for the one that was smaller than -3.. also do you believe that you can help me on another one of these because i can use the help i can get..

OpenStudy (the_fizicx99):

Yes B is correct! Sure!

OpenStudy (anonymous):

YES! and alright which point is NOT a solution of x+2y≤ 8? a. (0,4) b.(-2,3) c.(4,4) d. (6,0) i didnt really understand how he/she explained it to me..

OpenStudy (ipwnbunnies):

I thought you said you understood what I told you lol.

OpenStudy (ipwnbunnies):

Plug in the points, see if they satisfy or don't satisfy the inequality.

OpenStudy (anonymous):

i kinda did but i got lost in a few spots Cx lol sorry @iPwnBunnies

OpenStudy (anonymous):

alright let me see @iPwnBunnies

OpenStudy (anonymous):

Which ordered pair is NOT a solution of the inequality, 3x - 2y is greater than or equal to. 12 A. (-3, -10.5) B. (-1, -7) C. (0, -6) D. (2, -3) @tHe_FiZiCx99 wrong equation sorry

OpenStudy (the_fizicx99):

:< I should have added an explanation to my LaTeX :/ Ok, take the first option, A. \(\ \sf \Large (\color{blue}{X},\color{red}{Y}) \) So on option A, \(\ \rightarrow 0 \leftarrow \) is \(\ \sf \Large \color{blue}{X}\) And, \(\ \rightarrow 4 \leftarrow \) is \(\ \sf \Large \color{red}{Y}\) So, \(\ \sf \Large \color{blue}{0} + 2(\color{red}{4}) \le 8 \) \(\ \sf \Large \color{blue}{0} + \color{red}{8} \le 8 \longrightarrow 0 + 8 = 8 \) Now think, is \(\ \sf \Large 8 \le 8 \) ?

OpenStudy (the_fizicx99):

Oh there you made a typo X_X I'm going to take a break from using LaTeX o_o it's tiring >.< You could honestly plot this into a graphing calculator and see which ordered pair is not a solution to 3x + 2y ≥ 12

OpenStudy (anonymous):

no it isn't that one because its equal which mean its the right one

OpenStudy (the_fizicx99):

What?

OpenStudy (anonymous):

@tHe_FiZiCx99 ohhh i seee cx

OpenStudy (the_fizicx99):

:")

OpenStudy (anonymous):

sorry idk what i wrote Cx

OpenStudy (the_fizicx99):

Ok, so can you answer 3x + 2y ≥ 12 now on your own? If you write your steps I can tell you if you made any mistakes :>

OpenStudy (anonymous):

yeah i answered it , i was taking a quiz thats why it was online, and i still kjinda need help but its on percentages.

OpenStudy (the_fizicx99):

I'd suggest making a new thread and asking it there :P

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks alot :)

OpenStudy (the_fizicx99):

You're welcome!

OpenStudy (anonymous):

your good at math :)

OpenStudy (the_fizicx99):

Ty

OpenStudy (anonymous):

welcome :)